# Cointegration with Occasionally Binding Constraints

### v3

@@ 4.2 Case (ii): kinked cointegration
We turn next to the case in which the cointegrating rank $r$ is the same across both the positive and negative regimes, though the cointegrating space itself need not be. Here we also suppose that $\operatorname{rk}\Pi^{x}=r$, which as discussed in 3.3 entails that $y_{t}\sim I^{\ast}(1)$. Under the foregoing, we must have $\pi^{\pm}\in\operatorname{sp}\Pi^{x}$, and so
@@
Observe how (LABEL:pikink) implies that the ‘loadings’ $\alpha$ of the equilibrium errors will be the same in both regimes, even though the cointegrating vectors that define those errors need not be. (Case (iii) entails the opposite, with fixed cointegrating vectors but loadings that depend on the sign of $y_{t}$: see (LABEL:costatfactor) below.)
@@
[…]
@@
[…]
@@
(i). Similarly to 4.14.1 above, if DGP is replaced by DGP, then the theorem continues to hold exactly as stated, except that CO(ii).2 should be modified to
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via 2.1. (See D.)
(ii). Even when $\beta^{+}=\beta^{-}$, such that the cointegrating space is the same in both the positive and negative regimes, $(Y,X)$ will generally be a kinked Brownian motion because of the residual dependence of $P_{\beta_{\perp}}(y)=\beta_{\perp}[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}]^{-1}\alpha_{\perp}^{\mathsf{T}}$ on $y$ via $\Gamma(1;y)$. Indeed, in the univariate model (LABEL:univariate-case) under case (ii) with $r=0$,
@@
4.2 entails $Y(\lambda)=\gamma[1;U_{0}(\lambda)]^{-1}U_{0}(\lambda)$ is a kinked Brownian motion, whose variance depends on the sign of $U_{0}(\lambda)$.
(iii). CO(ii).2 plays an analogous role to CO(i).2 above, ensuring that the nonlinear VAR representation (see B.2) obtained for the short-memory components $(\xi_{t},\Delta z_{t})$ is sufficiently well-behaved that these processes are $I^{\ast}(0)$. Part (ii) would continue to hold if CO(ii).2 were replaced by any other condition sufficient for $(\xi_{t},\Delta z_{t})\sim I^{\ast}(0)$. A necessary condition for CO(ii).2 is that the eigenvalues of $I+\boldsymbol{\beta}(\pm 1)^{\mathsf{T}}\boldsymbol{\alpha}$ should lie strictly inside the unit circle, which is implied by CVAR (see A.3).
(iv). The map $(y,u)\mapsto P_{\beta_{\perp}}(y)u$ is not in general continuous, a fact that could interfere with the convergence in (LABEL:coclasXY). CO(ii).3 helps to ensure this map is continuous on a sufficiently large domain to permit (LABEL:coclasXY) to follow via an application of the continuous mapping theorem (CMT), with $Y$ and $X$ having continuous paths.
(v). Given $\alpha,\beta\in\mathbb{R}^{p\times r}$ with full column rank, their orthocomplements $\alpha_{\perp},\beta_{\perp}\in\mathbb{R}^{p\times q}$ are unique only up to their column span. Since in a linearly cointegrated system CVAR implies that $\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ has nonzero determinant (A.3), we may normalise $\alpha_{\perp}$ and/or $\beta_{\perp}$ so that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}=1$. It should therefore be emphasised that CO(ii).3 applies when $\beta_{\perp}(+1)$ and $\beta_{\perp}(-1)$ are related via (LABEL:betaperpfn), so we are not entirely free to choose $\beta_{\perp}(\pm 1)$ such that the signs of these determinants can be brought into agreement.
@@
To illustrate how 4.2 may be applied to derive the long-run behaviour of a structural model – even one in which the observables do not follow a (nonlinear) VAR – here we consider a slightly simplified version of the model of Ascari et al. (2023). They regard a vector of observable series $w_{t}$ as fluctuating in a stationary manner around their long-run components $\bar{w}_{t}$, as per
@@
where $A(L)$ is the lag polynomial of a stationary VAR. The long-run behaviour of $w_{t}$ will thus be governed by that of $\bar{w}_{t}$, which in our simplified formulation consists of trend inflation $\bar{\pi}_{t}$ and potential output $\bar{y}_{t}$. [^12] Their model also describes the evolution of the long-run component of the nominal interest rate; but since this variable does not feed back into the equations for $\bar{\pi}_{t}$ and $\bar{y}_{t}$, we may analyse the long-run behaviour of these processes separately from it. These are assumed, together with the trend growth rate $g_{t}$ of potential output, to evolve according to the nonlinear VAR
$\bar{\pi}_{t}$ $=\bar{\pi}_{t-1}+u_{t}^{\pi}$
$\Delta\bar{y}_{t}-\delta(\bar{\pi}_{t}-\tau)$ $=-\delta(\bar{\pi}_{t-1}-\tau)+g_{t-1}+u_{t}^{y}$
$g_{t}$ $=g_{t-1}+u_{t}^{g},$
where $\delta(\cdot)$ is a piecewise linear function of the form $\delta(x)=\delta^{+}[x]_{-}+\delta^{-}[x]_{+}$, which captures their ‘long-run Phillips curve’; see their equations (5)–(9). (Since we can trivially rewrite the model in terms of $\bar{\pi}_{t}-\tau$ rather than $\bar{\pi}_{t}$, here we also suppose that $\tau=0$, for simplicity.)
Because $\bar{\pi}_{t}$ enters this system nonlinearly, the long-run properties of these series cannot be determined on the basis of the ordinary Granger–Johansen representation theorem, but instead requires an application of the theory developed in the present work. To that end, we render the system in CKSVAR form as
@@
Direct calculation (for further details on this and the subsequent calculations, see F) yields
@@
so that $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=\operatorname{rk}\Pi^{x}=1$, indicating that the system falls within the purview of case (ii) with $\alpha=(0,1,0)^{\mathsf{T}}$ and $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$. It remains to verify the conditions of 4.2. We have immediately that CO(ii).1 holds with $r=1$. By calculating $\tilde{\Pi}^{\pm}$ for the associated canonical form of the model, we obtain that $01+\tilde{\beta}^{+\mathsf{T}}\tilde{\alpha}=1+\tilde{\beta}^{-\mathsf{T}}\tilde{\alpha}=$ 0, so (ii) is trivially satisfied. Finally, since in this first-order model $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, it may be shown that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma^{\pm}(1)\beta_{\perp}^{\pm}=\det I_{2}=1$, and thus CO(ii).3 holds. Since
@@
is regime-invariant, it follows therefore by 4.2 that (supposing that all processes are initialised at zero, for simplicity)
@@
where $n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}(u_{t}^{\pi},u_{t}^{g})\rightsquigarrow[U^{\pi}(\lambda),U^{g}(\lambda)]$. There are thus two common stochastic trends in the system, one of which is shared between $\Delta y_{t}$ and $g_{t}$; there is a single (linear) cointegrating relation given by the vector $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$, which eliminates these trends.
@@ 4.3 Case (iii): linear cointegration in a nonlinear VECM
where $\alpha_{x}\in\mathbb{R}^{p\times(r-1)}$ and $\beta_{x}\in\mathbb{R}^{(p-1)\times(r-1)}$ have rank $r-1$, and $\alpha^{\pm},\beta\in\mathbb{R}^{p\times r}$ have rank $r$. (Note that the dimensions of $\beta_{x}$ here differ from those in case (ii).) The model is thus one in which the positive and negative regimes share a common cointegrating space, which contains $e_{p,1}$. Thus we would expect $y_{t}$ to behave like a stationary process: and indeed, the model falls within the very general framework of Saikkonen (2008), whose results could be applied to establish the stationarity and ergodicity of the equilibrium errors
@@
and hence of $y_{t}$. Our technical contribution here is to exploit the structure of the CKSVAR so as to permit his conditions, which refer to the JSR of a collection of autoregressive matrices, to be relaxed to merely requiring the stability of a certain deterministic subsystem (for which control over the JSR is a sufficient but not necessary condition). [^13] It should be emphasised that we do not claim to be relaxing the conditions of Saikkonen (2008, Thm. 1) for the general class of regime-switching error correction models considered in that paper. Rather, we are able to exploit the fact that our model (the CKSVAR configured as per (LABEL:costatfactor)) is a special case of that framework to obtain weaker sufficient conditions for the ergodicity of $\beta^{\mathsf{T}}z_{t}$, in our setting.
@@
for $i\in\{1,\ldots,k\}$, $\boldsymbol{y}_{t-1}\coloneqq(y_{t-1},\ldots,y_{t-k})^{\mathsf{T}}$, and recognise that the factorisation (LABEL:costatfactor) applies more generally to
@@
for $\beta_{x,\perp}\in\mathbb{R}^{p-1}$ having $\operatorname{rk}\beta_{x,\perp}=p-r$, such that $\beta_{x,\perp}^{\mathsf{T}}\beta_{x}=0$. Letting $\boldsymbol{\chi}_{t}\coloneqq(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, we show in the proof of 4.3 below that
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via 2.1. The theorem then delivers the ${\cal Q}$-geometric ergodicity of $\tilde{\boldsymbol{\chi}}_{t}=(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, where
@@
is formed from the canonical parameters and variables. Since, as shown in D, we can write $\beta^{\mathsf{T}}z_{t}=(y_{t},(\beta_{x}^{\mathsf{T}}x_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\Delta x_{t}$ as measurable (indeed, Lipschitz continuous) functions of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$ and $(\tilde{\chi}_{t},\tilde{\chi}_{t-1})$ respectively, these processes will inherit the stationarity and geometric $\beta$-mixing properties of $\boldsymbol{\chi}_{t}$ that are a corollary of $\mathcal{Q}$-geometric ergodicity.
(ii). Since $\{\Delta x_{t}\}$ is geometrically $\beta$-mixing, and $x_{t}=\sum_{s=1}^{t}\Delta x_{s}+x_{0}$, the preceding result may be used as a starting point for the derivation of the asymptotics of $n^{-1/2}x_{\lfloor n\lambda\rfloor}$, which should converge to a multivariate Brownian motion (possibly after linear detrending). However, as discussed in Saikkonen (2008, pp. 307f.), determining the rank of the long-run variance of $\{\Delta x_{t}\}$, and hence the rank of the limiting process, is non-trivial. While its rank is bounded above by $q=p-r$, it need not be equal to $q$; guaranteeing the latter is likely to require further conditions on the model parameters. We leave this for future work.

### v4

@@ 4.2 Case (ii): kinked cointegration
We turn next to the case in which the cointegrating rank $r$ is the same across both the positive and negative regimes, though the cointegrating space itself need not be. Here we also suppose that $\operatorname{rk}\Pi^{x}=r$, which as discussed in Section 3.3 entails that $y_{t}\sim I^{\ast}(1)$. Under the foregoing, we must have $\pi^{\pm}\in\operatorname{sp}\Pi^{x}$, and so
@@
Observe how (4.6) implies that the ‘loadings’ $\alpha$ of the equilibrium errors will be the same in both regimes, even though the cointegrating vectors that define those errors need not be. (Case (iii) entails the opposite, with fixed cointegrating vectors but loadings that depend on the sign of $y_{t}$: see (4.15) below.)
@@
$\begin{aligned} \beta_{\perp}(y)&\coloneqq\begin{bmatrix}1&0\\ -\theta(y)&\beta_{x,\perp}\end{bmatrix},&\qquad\Gamma(1;y)&\coloneqq\Gamma^{+}(1)\mathbf{1}^{+}(y)+\Gamma^{-}(1)\mathbf{1}^{-}(y)\end{aligned},$ (4.8)
@@
      $n^{-1/2}z_{0}\overset{p}{\rightarrow}{\cal Z}_{0}=[\begin{smallmatrix}{\cal Y}_{0}\\ {\cal X}_{0}\end{smallmatrix}]\in\mathscr{M}$, where ${\cal Z}_{0}$ is non-random.
@@
(i). Similarly to Remark 4.14.1 above, if DGP is replaced by DGP, then the theorem continues to hold exactly as stated, except that CO(ii).2 should be modified to
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. (See Appendix D.)
(ii). Even when $\beta^{+}=\beta^{-}$, such that the cointegrating space is the same in both the positive and negative regimes, $(Y,X)$ will generally be a kinked Brownian motion because of the residual dependence of $P_{\beta_{\perp}}(y)=\beta_{\perp}[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}]^{-1}\alpha_{\perp}^{\mathsf{T}}$ on $y$ via $\Gamma(1;y)$. Indeed, in the univariate model (2.10) under case (ii) with $r=0$,
@@
Theorem 4.2 entails $Y(\lambda)=\gamma[1;U_{0}(\lambda)]^{-1}U_{0}(\lambda)$ is a kinked Brownian motion, whose variance depends on the sign of $U_{0}(\lambda)$.
(iii). CO(ii).2 plays an analogous role to CO(i).2 above, ensuring that the nonlinear VAR representation (see Lemma B.2) obtained for the short-memory components $(\xi_{t},\Delta z_{t})$ is sufficiently well-behaved that these processes are $I^{\ast}(0)$. Part (ii) would continue to hold if CO(ii).2 were replaced by any other condition sufficient for $(\xi_{t},\Delta z_{t})\sim I^{\ast}(0)$. A necessary condition for CO(ii).2 is that the eigenvalues of $I+\boldsymbol{\beta}(\pm 1)^{\mathsf{T}}\boldsymbol{\alpha}$ should lie strictly inside the unit circle, which is implied by CVAR (see Lemma A.3).
(iv). The map $(y,u)\mapsto P_{\beta_{\perp}}(y)u$ is not in general continuous, a fact that could interfere with the convergence in (4.11). CO(ii).3 helps to ensure this map is continuous on a sufficiently large domain to permit (4.11) to follow via an application of the continuous mapping theorem (CMT), with $Y$ and $X$ having continuous paths.
(v). Given $\alpha,\beta\in\mathbb{R}^{p\times r}$ with full column rank, their orthocomplements $\alpha_{\perp},\beta_{\perp}\in\mathbb{R}^{p\times q}$ are unique only up to their column span. Since in a linearly cointegrated system CVAR implies that $\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ has nonzero determinant (Lemma A.3), we may normalise $\alpha_{\perp}$ and/or $\beta_{\perp}$ so that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}=1$. It should therefore be emphasised that CO(ii).3 applies when $\beta_{\perp}(+1)$ and $\beta_{\perp}(-1)$ are related via (4.8), so we are not entirely free to choose $\beta_{\perp}(\pm 1)$ such that the signs of these determinants can be brought into agreement.
@@
To illustrate how Theorem 4.2 may be applied to derive the long-run behaviour of a structural model – even one in which the observables do not follow a (nonlinear) VAR – here we consider the model of Ascari et al. (2023). They regard a vector of observable series $w_{t}$ (the inflation rate, GDP per capita, and the nominal interest rate) as fluctuating in a stationary manner around their long-run components $\bar{w}_{t}$, as per
@@
where $A(L)$ is the lag polynomial of a stationary VAR. The long-run behaviour of $w_{t}$ will thus be governed by that of $\bar{w}_{t}$, which consists of trend inflation $\bar{\pi}_{t}$, potential output $\bar{y}_{t}$, and the trend nominal interest rate $\bar{\imath}_{t}$. The first two components are assumed, together with the trend growth rate $g_{t}$ of potential output, to evolve according to the nonlinear VAR
$\bar{\pi}_{t}$ $=\bar{\pi}_{t-1}+u_{t}^{\pi}$ (4.12a)
$\Delta\bar{y}_{t}-\delta(\bar{\pi}_{t}-\tau)$ $=-\delta(\bar{\pi}_{t-1}-\tau)+g_{t-1}+u_{t}^{y}$ (4.12b)
$g_{t}$ $=g_{t-1}+u_{t}^{g},$ (4.12c)
where $\delta(\cdot)$ is a piecewise linear function of the form $\delta(x)=\delta^{+}[x]_{-}+\delta^{-}[x]_{+}$, which captures their ‘long-run Phillips curve’; see their equations (5)–(9). (Since we can trivially rewrite the model in terms of $\bar{\pi}_{t}-\tau$ rather than $\bar{\pi}_{t}$, here we also suppose that $\tau=0$, for simplicity.) The trend in the nominal interest rate is specified to be a linear function of $\bar{\pi}_{t}$, $g_{t}$, and an additional stochastic trend $z_{t}$, as per their equations (10) and (11):
$\bar{\imath}_{t}$ $=\bar{\pi}_{t}+cg_{t}+z_{t}$ (4.13a)
$z_{t}$ $=z_{t-1}+u_{t}^{z}.$ (4.13b)
Since $\bar{\imath}_{t}$ does not enter (4.12), we can analyse the long-run behaviour of this subsystem separately, and thence deduce that of $\bar{\imath}_{t}$.
Because $\bar{\pi}_{t}$ enters (4.12) nonlinearly, the long-run properties of $(\bar{\pi},\Delta\bar{y}_{t},g_{t})$ cannot be determined on the basis of the ordinary Granger–Johansen representation theorem, but instead requires an application of the theory developed in the present work. To that end, we render the system (4.12) in CKSVAR form as
@@
Direct calculation (for further details on this and the subsequent calculations, see Appendix F) yields
@@
so that $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=\operatorname{rk}\Pi^{x}=1$, indicating that the system falls within the purview of case (ii) with $\alpha=(0,1,0)^{\mathsf{T}}$ and $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$. It remains to verify the conditions of Theorem 4.2. We have immediately that CO(ii).1 holds with $r=1$. By calculating $\tilde{\Pi}^{\pm}$ for the associated canonical form of the model, we obtain that $1+\tilde{\beta}^{+\mathsf{T}}\tilde{\alpha}=1+\tilde{\beta}^{-\mathsf{T}}\tilde{\alpha}=0$, so (ii) is trivially satisfied. Finally, since in this first-order model $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, it may be shown that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma^{\pm}(1)\beta_{\perp}^{\pm}=\det I_{2}=1$, and thus CO(ii).3 holds. Since
@@
is regime-invariant, it follows therefore by Theorem 4.2 that (supposing that all processes are initialised at zero, for simplicity)
@@
where $n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}(u_{t}^{\pi},u_{t}^{g})\rightsquigarrow[U^{\pi}(\lambda),U^{g}(\lambda)]$.
There are thus two common stochastic trends in the subsystem (4.12), one of which is shared between $\Delta\bar{y}_{t}$ and $g_{t}$; there is a single (linear) cointegrating relation given by the vector $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$, which eliminates these trends. It follows moreover from (4.13) that
$n^{-1/2}\bar{\imath}_{\lfloor n\lambda\rfloor}=n^{-1/2}(\bar{\pi}_{\lfloor n\lambda\rfloor}+cg_{\lfloor n\lambda\rfloor}+z_{\lfloor n\lambda\rfloor})\rightsquigarrow U^{\pi}(\lambda)+cU^{g}(\lambda)+U^{z}(\lambda),$
where $n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}u_{t}^{z}\rightsquigarrow U^{z}(\lambda)$, and so $\bar{\imath}_{t}$ does not cointegrate with $(\bar{\pi},\Delta\bar{y}_{t},g_{t})$.
@@ 4.3 Case (iii): linear cointegration in a nonlinear VECM
where $\alpha_{x}\in\mathbb{R}^{p\times(r-1)}$ and $\beta_{x}\in\mathbb{R}^{(p-1)\times(r-1)}$ have rank r1r-1, and $\alpha^{\pm},\beta\in\mathbb{R}^{p\times r}$ have rank $r$. (Note that the dimensions of $\beta_{x}$ here differ from those in case (ii).) The model is thus one in which the positive and negative regimes share a common cointegrating space, which contains $e_{p,1}$. Thus we would expect $y_{t}$ to behave like a stationary process: and indeed, the model falls within the very general framework of Saikkonen (2008), whose results could be applied to establish the stationarity and ergodicity of the equilibrium errors
@@
and hence of $y_{t}$. Our technical contribution here is to exploit the structure of the CKSVAR so as to permit his conditions, which refer to the JSR of a collection of autoregressive matrices, to be relaxed to merely requiring the stability of a certain deterministic subsystem (for which control over the JSR is a sufficient but not necessary condition). [^12] It should be emphasised that we do not claim to be relaxing the conditions of Saikkonen (2008, Thm. 1) for the general class of regime-switching error correction models considered in that paper. Rather, we are able to exploit the fact that our model (the CKSVAR configured as per (4.15)) is a special case of that framework to obtain weaker sufficient conditions for the ergodicity of $\beta^{\mathsf{T}}z_{t}$, in our setting.
@@
for $i\in\{1,\ldots,k\}$, $\boldsymbol{y}_{t-1}\coloneqq(y_{t-1},\ldots,y_{t-k})^{\mathsf{T}}$, and recognise that the factorisation (4.15) applies more generally to
@@
for $\beta_{x,\perp}\in\mathbb{R}^{p-1}$ having $\operatorname{rk}\beta_{x,\perp}=p-r$, such that $\beta_{x,\perp}^{\mathsf{T}}\beta_{x}=0$. Letting $\boldsymbol{\chi}_{t}\coloneqq(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, we show in the proof of Theorem 4.3 below that
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. The theorem then delivers the $\cal Q$}-geometric ergodicity of $\tilde{\boldsymbol{\chi}}_{t}=(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, where
@@
is formed from the canonical parameters and variables. Since, as shown in Appendix D, we can write $\beta^{\mathsf{T}}z_{t}=(y_{t},(\beta_{x}^{\mathsf{T}}x_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\Delta x_{t}$ as measurable (indeed, Lipschitz continuous) functions of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$ and $(\tilde{\chi}_{t},\tilde{\chi}_{t-1})$ respectively, these processes will inherit the stationarity and geometric $\beta$-mixing properties of $\boldsymbol{\chi}_{t}$ that are a corollary of $\mathcal{Q}$-geometric ergodicity.
(ii). Since $\{\Delta x_{t}\}$ is geometrically $\beta$-mixing, and $x_{t}=\sum_{s=1}^{t}\Delta x_{s}+x_{0}$, the preceding result may be used as a starting point for the derivation of the asymptotics of $n^{-1/2}x_{\lfloor n\lambda\rfloor}$, which should converge to a multivariate Brownian motion (possibly after linear detrending). However, as discussed in Saikkonen (2008, pp. 307f.), determining the rank of the long-run variance of $\{\Delta x_{t}\}$, and hence the rank of the limiting process, is non-trivial. While its rank is bounded above by q=prq=p-r, it need not be equal to $q$; guaranteeing the latter is likely to require further conditions on the model parameters. We leave this for future work.